I need help with finding the inverse of:
f(x)-2x+7

Answer by tutorcecilia(2152)

f(x)= -2x+7
x=-2y+7 [Switch x an y]
x-7=-2y [Solve for y]
(x-7)/-2=y
So, f^-1(x) = (x-7)/-2
.
Checking by using (f of f^-1)=x
f(x)= -2x+7
(f of f^-1) = -2((x-7)/-2)+7
(x-7)+7
x
.
Also, checking by using (f^-1 of f)=x
(f^-1 of f) = ((-2x+7)-7)/-2
-2x/-2
x [Both functions check out]